OpenStudy (anonymous):

Please answer !!! need help :) Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (anonymous):

They have one factor in common..

OpenStudy (anonymous):

\[x^2 + 3x + 2 = (x+1)(x+2)\] \[x^2 + 2x + 1 = (x+1)(x+1)\] So, \((x+1)\) is common in both..

OpenStudy (anonymous):

There b and c values are different which is the sufficient condition to make them different..

OpenStudy (anonymous):

thank you